# Complexity of Local Search for Euclidean Clustering Problems

### v1

@@ Abstract
We show that the simplest local search heuristics for two natural Euclidean clustering problems are PLS-complete. First, we show that the Hartigan–Wong method for $k$-Means clustering is PLS-complete, even when $k=2$. Second, we show the same result for the Flip heuristic for Max Cut, even when the edge weights are given by the (squared) Euclidean distances between the points in some set $\mathcal{X}\subseteq\mathds{R}^{d}$; a problem which is equivalent to Min Sum 2-Clustering.
keywords:
Local search, PLS-complete, max cut, k-means, partitioning problem, flip-neighborhood
@@ 1 Introduction
$k$-Means
One well-studied clustering problem is $k$-Means [berkhinSurveyClusteringData2006, jainDataClustering502010]. In this problem, one is given a set of points $\mathcal{X}\subseteq\mathds{R}^{d}$ and an integer $k$. The goal is to partition $\mathcal{X}$ into exactly $k$ clusters such that the total squared distance of each point to the centroid of its cluster is minimized. Formally, one seeks to minimize the clustering cost
@@
Being NP-hard even when $k=2$ [ADHP09] or when $\mathcal{X}\subseteq\mathds{R}^{2}$ [mahajanPlanarKmeansProblem2012], $k$-Means has been extensively studied from the perspective of approximation algorithms, with several such approaches having emerged [arthurKmeansAdvantagesCareful2007, hasegawaEfficientAlgorithmsVarianceBased2000, kanungoEfficientKmeansClustering2002, matousekApproximateGeometricClustering2000]. Nevertheless, local search remains the method of choice for practitioners [berkhinSurveyClusteringData2006, jainDataClustering502010].
@@
The most well-known local search algorithm for $k$-Means is Lloyd’s method [lloydLeastSquaresQuantization1982]. Here, one alternates between two steps in each iteration. In the first step, each point is assigned to its closest cluster center, and in the second step the cluster centers are recalculated from the newly formed clusters.
@@
This algorithm was shown to have worst-case super-polynomial running time by Arthur and Vassilvitskii [arthurHowSlowKmeans2006], with the result later improved to exponential running time even in the plane by Vattani [vattaniKmeansRequiresExponentially2011]. Moreover, Roughgarden and Wang [roughgardenComplexityKmeansMethod2016] showed it can implicitly solve PSPACE-complete problems. On the other hand, Arthur et al. [arthurSmoothedAnalysisKMeans2011] proved that Lloyd’s method has smoothed polynomial running time on Gaussian-perturbed point sets, providing a degree of explanation for its effectiveness in practice.
@@
Recently Telgarsky and Vattani [telgarskyHartiganMethodKmeans2010] revived interest in another, older local search method for $k$-Means due to Hartigan and Wong [hartiganAlgorithm136KMeans1979]. This algorithm, the Hartigan–Wong method, is much simpler: one searches for a single point that can be reassigned to some other cluster for a strict improvement in the clustering cost. Despite this simplicity, they show that the Hartigan–Wong method is more powerful than Lloyd’s method, in the sense that the former can sometimes improve clusterings produced by the latter, while the converse does not hold.
@@
A similar construction to that of Vattani for Lloyd’s method shows that there exist instances on which the Hartigan–Wong method can take exponentially many iterations to find a local optimum, even when all points lie on a line [mantheyWorstCaseSmoothedAnalysis2023]. However, this example follows a contrived sequence of iterations. Moreover, $k$-Means can be solved optimally for instances in which all points lie on a line. Thus, the question remains whether stronger worst-case examples exist, and what the complexity of finding locally optimal clusterings is.
Squared Euclidean Max Cut
@@
Another clustering problem similar to $k$-Means is Squared Euclidean Max Cut. Recall that Max Cut asks for a subset of vertices $S$ of a weighted graph $G=(V,E)$, such that the total weight of the edges with one endpoint in $S$ and one in $V\setminus S$ is maximized. This problem emerges in numerous applications, from graph clustering to circuit design to statistical physics [barahonaApplicationCombinatorialOptimization1988, boykovInteractiveGraphCuts2001].
@@
over all partitions $(X,Y)$ of $\mathcal{X}$. Also this special case of Max Cut is NP-hard [ageevComplexityWeightedMaxcut2014]. In a clustering context, the problem was studied by Schulman [schulmanClusteringEdgecostMinimization2000] and Hasegawa et al. [hasegawaEfficientAlgorithmsVarianceBased2000], leading to exact and approximation algorithms.
@@
Given the computational hardness of Max Cut, practitioners often turn to heuristic solutions. Some of the resulting algorithms are very successful, such as the Kernighan-Lin heuristic [kernighanEfficientHeuristicProcedure1970] and the Fiduccia-Mattheyses algorithm [fiducciaLinearTimeHeuristicImproving1982a]. Johnson et al. [johnsonHowEasyLocal1988] note that the simple Flip heuristic, where one moves a single vertex from one side of the cut to the other, tends to converge quickly in practice (note the similarity with the Hartigan–Wong method). Schäffer and Yannakakis [schafferSimpleLocalSearch1991] later showed that it has exponential running time in the worst case.
@@
One may wonder whether Flip performs better for Squared Euclidean Max Cut. Etscheid and Röglin [etscheidSmoothedAnalysisSquared2015a, etscheidWorstCaseAnalysisMaxCut2018] performed a smoothed analysis of the Flip heuristic in this context, showing a smoothed running time of $2^{O(d)}\cdot\operatorname{poly}(n,1/\sigma)$ for Gaussian-perturbed instances. On the other hand, they also exhibited an instance in $\mathds{R}^{2}$ on which there exists an exponential-length improving sequence of iterations, with the caveat that not all edges are present in the instance. Like for $k$-Means, one may ask whether stronger examples exist (e.g. on complete graphs), and what is the complexity of finding Flip-optimal solutions.
Complexity of Local Search
@@
The existence of instances with worst-case exponential running time is common for local search heuristics. To investigate this phenomenon, and local search heuristics in general, Johnson et al. [johnsonHowEasyLocal1988] defined a complexity class PLS, for polynomial local search. The class is designed to capture the properties of commonly used local search heuristics and contains pairs consisting of an optimization problem P and a local search heuristic $\mathcal{N}$. In the following we denote such a pair as P/$\mathcal{N}$. Complete problems for PLS have the property that their natural local search algorithms have worst-case exponential running time. In the same work, Johnson et al. showed that the Kernighan-Lin heuristic for the Max Bisection problem (a variant of Max Cut, where the parts of the partition must be of equal size) is PLS-complete. This stands in contrast to the empirical success of this algorithm [kernighanEfficientHeuristicProcedure1970].
@@
Building on this work, Schäffer and Yannakakis [schafferSimpleLocalSearch1991] later proved that a host of very simple local search heuristics are PLS-complete, including the Flip heuristic for Max Cut. This refuted a conjecture by Johnson et al., who doubted that such simple heuristics could be PLS-complete. Elsässer and Tscheuschner [elsasserSettlingComplexityLocal2011] later showed that this remains true even in the very restricted variant where the input graph has maximum degree five, which we will refer to as Max Cut-5.
@@
Since the introduction of PLS, many local search problems have been shown to be PLS-complete, including such successful heuristics as Lin-Kernighan’s algorithm for the TSP [papadimitriouComplexityLinKernighan1992] or the $k$-Swap-neighborhood heuristic for Weighted Independent Set [komusiewiczMorawietzFindingSwap2022] for $k\geq 3$. For a non-exhaustive list, see Michiels, Korst and Aarts [michielsTheoreticalAspectsLocal2007, Appendix C].
Our Contribution
Given the existence of exponential worst-case instancs of $k$-Means/Hartigan–Wong, one may ask whether this heuristic is PLS-complete. In this work, we answer this question in the affirmative.
Theorem 1.1.
$k$-Means/Hartigan–Wong is PLS-complete, even if $k=2$.
Just as with $k$-Means/Hartigan–Wong, we ask whether Squared Euclidean Max Cut with the Flip heuristic is PLS-complete. Again, we answer this question affirmatively. In addition, we show the same result for Euclidean Max Cut, where the distances between the points are not squared.
Theorem 1.2.
@@
We note that PLS-completeness results for Euclidean local optimization problems are rather uncommon. We are only aware of one earlier result by Brauer [brauerComplexitySingleSwapHeuristics2017], who proved PLS-completeness of a local search heuristic for a discrete variant of $k$-Means. This variant chooses $k$ cluster centers among the set of input points, after which points are assigned to their closest center. The heuristic they consider removes one point from the set of cluster centers and adds another. In their approach, they first construct a metric instance, and then show that this instance can be embedded into $\mathds{R}^{d}$ using a theorem by Schoenberg [schoenbergMetricSpacesPositive1938]. In contrast, we directly construct instances in $\mathds{R}^{d}$.
In addition to showing PLS-hardness of $k$-Means/Hartigan–Wong and Squared Euclidean Max Cut/Flip, we also show that there exist specific hard instances of these problems, as well as all other problems considered in this paper.
Theorem 1.3.
@@
    There exists an instance of $L$ with an initial solution that is exponentially far away from any local optimum.
@@
    The problem of computing the locally optimal solution obtained from performing a standard hill-climbing algorithm is PSPACE-complete for $L$.
A formal definition of Property 3 is given in Section 2.1. In particular, this result shows that there exists an instance of $k$-Means such that there exists an initial solution to this instance that is exponentially many iterations removed from all local optima [schafferSimpleLocalSearch1991]. By contrast, the earlier result [mantheyWorstCaseSmoothedAnalysis2023] only exhibits an instance with a starting solution that is exponentially many iterations away from some local optimum. Moreover, Theorem 1.3 yields instances where Flip has exponential running time in Squared Euclidean Max Cut on complete geometric graphs, unlike the older construction which omitted some edges [etscheidWorstCaseAnalysisMaxCut2018].
@@ 2.1 The Class PLS
For convenience, we summarize the formal definitions of local search problems and the associated complexity class PLS, as devised by Johnson et al. [johnsonHowEasyLocal1988].
@@
The class PLS is defined to capture the properties of local search problems that typically arise in practical applications. Formally, $P$ is said to belong to PLS if the following are all true:
@@
    There exists a polynomial-time algorithm $A$ that, given an instance $x\in I$, computes some solution $s\in F_{I}(x)$.
@@
    There exists a polynomial-time algorithm $B$ that, given $x\in I$ and $s\in F_{I}(x)$, computes the value of $c(s,x)$.
@@
    There exists a polynomial-time algorithm $C$ that, given $x\in I$ and $s\in F_{I}(x)$, either computes a solution $s^{\prime}\in\mathcal{N}(s,x)$ with $c(s^{\prime},x)>c(s,x)$ (in the case of a maximization problem), or outputs that such a solution does not exist.
@@
Definition 2.2.
Let $P$ be a local search problem and let $I$ be an instance of P. Moreover, let $s^{*}(I)$ be the unique local optimum obtained by starting with the solution outputted by algorithm $A$ and replacing the current solution by the better solution outputted by $C$, until reaching a local optimum. The standard algorithm problem for $P$ asks for a given instance $I$ of $P$ and a locally optimal solution $s^{\prime}$ for $I$ with respect to $\mathcal{N}$, wether $s^{\prime}$ is exactly the solution $s^{*}(I)$.
It was shown that for many local search problems the standard algorithm problem is PSPACE-complete [brauerComplexitySingleSwapHeuristics2017, michielsTheoreticalAspectsLocal2007, monienPowerNodesDegree2010, schafferSimpleLocalSearch1991].
@@
    There exist polynomial-time computable functions $f$ and $g$, such that $f$ maps instances $x$ of $P$ to instances $f(x)$ of $Q$, and $g$ maps pairs (solution s of $(\text{solution $s$ of $f(x)$},x)$ to feasible solutions of $P$.
@@
In addition to this standard notion of a PLS-reduction, Schäffer and Yannakakis [schafferSimpleLocalSearch1991] defined so-called tight reductions. Given PLS problems $P$ and $Q$ and a PLS-reduction $(f,g)$ from $P$ to $Q$, the reduction is called tight if for any instance $x$ of $P$ we can choose a subset $\mathcal{R}$ of the feasible solutions of $f(x)$ of $Q$ such that:
@@
Tight PLS-reductions have two desired properties [aartsLocalSearchCombinatorial2003, Chapter 2]. Suppose $P$ reduces to $Q$ via a tight reduction. First, if the standard algorithm problem for $P$ is PSPACE-complete, then the standard algorithm problem for $Q$ is PSPACE-complete as well. Second, if there exists an instance $x$ of $P$ such that there exists a solution of $x$ that is exponentially far away from any local optimum, then such an instance exists for $Q$ as well.
@@ 2.2 Definitions of Local Search Problems
We will be concerned with various local search problems. In the following we provide a summary of the problems that appear in this paper, and provide definitions for each. Most of the problems considered in this paper are not the most natural ones, but we need them as intermediate problems for our reductions. Moreover, these problems might be useful to show PLS-hardness of other problems having cardinality constraints.
@@
The Flip neighborhood as defined above is perhaps the simplest neighborhood structure for a partioning problem. For each problem in the list below, we consider only the Flip neighborhood in this paper. For $k$-Means, the neighborhood is called Hartigan–Wong, which is synonymous with Flip; the name stems from the paper by Hartigan and Wong which introduced the neighborhood [hartiganAlgorithm136KMeans1979].
@@
Input:
A graph $G=(V,E)$ with non-negative edge weights $w:E\to\mathds{Z}_{\geq 0}$.
Output:
A partition $(X,Y)$ of $V$ such that $w(X,Y)$ is maximal.
@@
The definitions of Odd Max/Min Bisection are somewhat unconventional, as one usually considers these problem with an even number of vertices and with the Swap neighborhood, where two solutions are neighbors if one can be obtained from the other by swapping a pair of vertices between parts of the partition. Hardness of Max Bisection/Swap was shown by Schäffer and Yannakakis [schafferSimpleLocalSearch1991] in a short reduction from Max Cut/Flip, providing as a corollary also a simple hardness proof for the Kernighan-Lin heuristic [kernighanEfficientHeuristicProcedure1970]. The reason we require the Flip neighborhood is that we aim to reduce this problem to Squared Euclidean Max Cut/Flip, and we run into trouble when we use the Swap neighborhood.
@@
Input:
A complete graph $G=(V,E)$ with non-negative edge weights $w:E\to\mathds{Z}_{\geq 0}$, such that there exists an injection $f:V\to\mathds{R}^{d}$ such that $w(uv)=\|f(u)-f(v)\|^{2}$ for all edges $uv\in E$.
Output:
A partition $(X,Y)$ of $V$ such that $w(X,Y)$ is maximal.
Euclidean Max Cut is defined similarly; the only difference is that we have $w(uv)=\|f(u)-f(v)\|$.
Squared Euclidean Max Cut is more commonly defined in terms of an input point set $\mathcal{X}\subseteq\mathds{R}^{d}$, with solutions being partitions $(X,Y)$ of $\mathcal{X}$ with cost $\sum_{x\in X}\sum_{y\in Y}\|x-y\|^{2}$. However, in this formulation, membership of Squared Euclidean Max Cut/Flip in PLS is no longer necessarily true111 The problem Sum of Square Roots asks whether for given numbers $a_{i}$ and $k_{i}$ the linear combination $\sum_{i}k_{i}\sqrt{a_{i}}$ evaluates to zero (among other variants). It is not currently known whether this problem can be solved in time polynomial in the number of bits needed to represent the given linear combination, although a randomized algorithm can solve the problem in polynomial time [blomerComputingSumsRadicals1991]. Since computing the cost of a solution and deciding local optimality for (Squared) Euclidean Max Cut/Flip and $k$-Means/Hartigan–Wong requires the computation of such linear combinations when formulated in the traditional way, it is not clear whether these formulations are in PLS. . Hence, we opt for the above graph-theoretical formulation.
@@
Input:
A complete graph $G=(V,E)$ with non-negative edge weights $w:E\to\mathds{Z}_{\geq 0}$, such that there exists a injection $f:V\to\mathds{R}^{d}$ such that $w(uv)=\|f(u)-f(v)\|^{2}$ for all edges $uv\in E$.
Output:
A partition $(C_{1},\ldots,C_{k})$ of $V$ such that $\sum_{i=1}^{k}\frac{1}{|C_{i}|}\sum_{u,v\in C_{i}}w(uv)$ is minimal.
The sets $\{C_{1},\ldots,C_{k}\}$ are also referred to as clusters. As with Squared Euclidean Max Cut, in the literature $k$-Means is usually specified with an input point set $\mathcal{X}\subseteq\mathds{R}^{d}$, and the cost of a partition $\{C_{1},\ldots,C_{k}\}$ of $\mathcal{X}$ is given as $\sum_{i=1}^{k}\sum_{x\in C_{i}}\|x-\operatorname{cm}(C_{i})\|^{2}$, where $\operatorname{cm}(C_{i})$ denotes to centroid of $C_{i}$. This is equivalent to our formulation when the coordinates of the points are algebraic numbers [pyatkinNPHardnessBalancedMinimum2017, Proposition 1]. In LABEL:lemma:densestcut_to_hartigan and LABEL:lemma:k-means we will also use this notation.
Positive Not-All-Equal $k$-Satisfiability (Pos NEA $k$-SAT)
Input:
A boolean formula with clauses of the form $\operatorname{NAE}(x_{1},\ldots,x_{\ell})$ with $\ell\leq k$, where each clause is satisfied if its constituents, all of which are positive, are neither all true nor all false. Each clause $C$ has a weight $w(C)\in\mathds{Z}$.
Output:
A truth assignment of the variables such that the sum of the weights of the satisfied clauses is maximized.
In Odd Half Pos NEA $k$-SAT, additionally the number of variables is odd and it is required that the number of true variables and the number of false variables in any solution differ by exactly one. This is analogous to the relationship between Max Cut and Odd Max Bisection.

### v3

@@ Abstract
We show that the simplest local search heuristics for two natural Euclidean clustering problems are PLS-complete. First, we show that the Hartigan–Wong method, which is essentially the Flip heuristic, for $k$-Means clustering is PLS-complete, even when $k=2$. Second, we show the same result for the Flip heuristic for Max Cut, even when the edge weights are given by the (squared) Euclidean distances between the points in some set $\mathcal{X}\subseteq\mathds{R}^{d}$; a problem which is equivalent to Min Sum 2-Clustering.
@@ 1 Introduction
$k$-Means.
One well-studied clustering problem is $k$-Means [10, 24]. In this problem, one is given a set of points $\mathcal{X}\subseteq\mathds{R}^{d}$ and an integer $k$. The goal is to partition $\mathcal{X}$ into exactly $k$ clusters such that the total squared distance of each point to the centroid of its cluster is minimized. Formally, one seeks to minimize the clustering cost
@@
Being NP-hard even when $k=2$ [3] or when $\mathcal{X}\subseteq\mathds{R}^{2}$ [32], $k$-Means has been extensively studied from the perspective of approximation algorithms, with several such approaches having emerged[8, 22, 27, 35]. Nevertheless, local search remains the method of choice for practitioners [10, 24].
@@
The most well-known local search algorithm for $k$-Means is Lloyd’s method [31]. Here, one alternates between two steps in each iteration. In the first step, each point is assigned to its closest cluster center, and in the second step the cluster centers are recalculated from the newly formed clusters.
@@
This algorithm was shown to have worst-case super-polynomial running time by Arthur and Vassilvitskii [7], with the result later improved to exponential running time even in the plane by Vattani [47]. Moreover, Roughgarden and Wang [41] showed it can implicitly solve PSPACE-complete problems. On the other hand, Arthur et al. [6] proved that Lloyd’s method has smoothed polynomial running time on Gaussian-perturbed point sets, providing a degree of explanation for its effectiveness in practice.
@@
Recently Telgarsky and Vattani [46] revived interest in another, older local search method for $k$-Means due to Hartigan and Wong [21]. This algorithm, the Hartigan–Wong method, is much simpler: one searches for a single point that can be reassigned to some other cluster for a strict improvement in the clustering cost. In other words, the Hartigan–Wong method is the Flip heuristic. In the following, we always use Flip instead of Hartigan–Wong to indicate that this is the most simple heuristic for this problem and to keep the name for the used heuristic consistent. Despite this simplicity, Telgarsky and Vattani [46] show that the Flip heuristic is more powerful than Lloyd’s method, in the sense that the former can sometimes improve clusterings produced by the latter, while the converse does not hold.
@@
A similar construction to that of Vattani for Lloyd’s method shows that there exist instances on which the Flip heuristic can take exponentially many iterations to find a local optimum, even when all points lie on a line [34]. However, this example follows a contrived sequence of iterations. Moreover, $k$-Means can be solved optimally for instances in which all points lie on a line. Thus, the question remains whether stronger worst-case examples exist, and what the complexity of finding locally optimal clusterings is.
Squared Euclidean Max Cut.
@@
Another clustering problem similar to $k$-Means is Squared Euclidean Max Cut. Recall that Max Cut asks for a subset of vertices $S$ of a weighted graph $G=(V,E)$, such that the total weight of the edges with one endpoint in $S$ and one in $V\setminus S$ is maximized. This problem emerges in numerous applications, from graph clustering to circuit design to statistical physics [9, 13].
@@
over all partitions $(X,Y)$ of $\mathcal{X}$. Also this special case of Max Cut is NP-hard [2]. In a clustering context, the problem was studied by Schulman [44] and Hasegawa et al. [22], leading to exact and approximation algorithms.
@@
Given the computational hardness of Max Cut, practitioners often turn to heuristics. Some of the resulting algorithms are very successful, such as the Kernighan-Lin heuristic [28] and the Fiduccia-Mattheyses algorithm [20]. Johnson et al. [25] note that the simple Flip heuristic, where one moves a single vertex from one side of the cut to the other, tends to converge quickly in practice. Schäffer and Yannakakis [42] later showed that it has exponential running time in the worst case.
@@
One may wonder whether Flip performs better for Squared Euclidean Max Cut. Etscheid and Röglin [18, 17] performed a smoothed analysis of Flip in this context, showing a smoothed running time of $2^{O(d)}\cdot\operatorname{poly}(n,1/\sigma)$ for Gaussian-perturbed instances, where $\sigma$ denotes the standard deviation of the Gaussian noise. On the other hand, they also exhibited an instance in $\mathds{R}^{2}$ on which there exists an exponential-length improving sequence of iterations, with the caveat that not all edges are present in the instance. Like for $k$-Means, one may ask whether stronger examples exist (e.g. on complete graphs), and what the complexity of finding Flip-optimal solutions is.
Complexity of Local Search.
@@
The existence of instances with worst-case exponential running time is common for local search heuristics. To investigate this phenomenon, and local search heuristics in general, Johnson et al. [25] defined a complexity class PLS, for polynomial local search. The class is designed to capture the properties of commonly used local search heuristics and contains pairs consisting of an optimization problem P and a local search heuristic $\mathcal{N}$. In the following we denote such a pair as P/$\mathcal{N}$. PLS-complete problems have the property that their natural local search algorithms have worst-case exponential running time. In the same work, Johnson et al. showed that the Kernighan-Lin heuristic for the Max Bisection problem (a variant of Max Cut, where the parts of the partition must be of equal size) is PLS-complete. This stands in contrast to the empirical success of this algorithm [28].
@@
Building on this work, Schäffer and Yannakakis [42] later proved that a host of very simple local search heuristics are PLS-complete, including the Flip heuristic for Max Cut. This refuted a conjecture by Johnson et al., who doubted that such simple heuristics could be PLS-complete. Elsässer and Tscheuschner [16] later showed that this remains true even in the very restricted variant where the input graph has maximum degree five, which we will refer to as Max Cut-5.
@@
Since the introduction of PLS, many local search problems have been shown to be PLS-complete, including such successful heuristics as Lin-Kernighan’s algorithm for the TSP [39] or the $k$-Swap-neighborhood heuristic for Weighted Independent Set [29] for $k\geq 3$. For a non-exhaustive list, see Michiels, Korst and Aarts [37, Appendix C].
Our Contribution.
Given the existence of $k$-Means instances where the Flip heuristic has worst-case exponential running time, one may ask whether this heuristic is PLS-complete. In this work, we answer this question in the affirmative.
Theorem 1.
For each $k\geq 2$, $k$-Means/Flip is PLS-complete.
Just as with $k$-Means/Flip, we ask whether Squared Euclidean Max Cut with the Flip heuristic is PLS-complete. Again, we answer this question affirmatively. In addition, we show the same result for Euclidean Max Cut, where the distances between the points are not squared.
Theorem 2.
@@
We note that PLS-completeness results for Euclidean local optimization problems are rather uncommon. We are only aware of one earlier result by Brauer [14], who proved PLS-completeness of a local search heuristic for a discrete variant of $k$-Means. This variant chooses $k$ cluster centers among the set of input points, after which points are assigned to their closest center. The heuristic they consider removes one point from the set of cluster centers and adds another. In their approach, they first construct a metric instance, and then show that this instance can be embedded into $\mathds{R}^{d}$ using a theorem by Schoenberg [43]. In contrast, we directly construct instances in $\mathds{R}^{d}$.
In addition to showing PLS-hardness of $k$-Means/Flip and Squared Euclidean Max Cut/Flip, we also show that there exist specific hard instances of these problems, as well as all other problems considered in this paper.
Theorem 3.
@@
    For each $n$, one can compute in polynomial time an instance of $L$ of size $n^{\mathcal{O}(1)}$ with an initial solution that is exponentially far away from any local optimum.
@@
    The problem of computing the locally optimal solution obtained from performing a standard local search algorithm based on the neighborhood of $L$ is PSPACE-complete for $L$.
A formal definition of Property 3 is given in Section 2.1. In particular, this result shows that there exists an instance of $k$-Means such that there exists an initial solution to this instance that is exponentially many iterations away from all local optima [42]. By contrast, the earlier result [34] only exhibits an instance with a starting solution that is exponentially many iterations away from some local optimum. Moreover, Theorem 3 yields instances where Flip has exponential running time in Squared Euclidean Max Cut on complete geometric graphs, unlike the older construction which omitted some edges [17].
@@ 2.1 The Class PLS
For convenience, we summarize the formal definitions of local search problems and the associated complexity class PLS, as devised by Johnson et al. [25].
@@
The class PLS is defined to capture the properties of local search problems that typically arise in practical applications. Formally, $P$ is contained in PLS if the following are all true:
@@
    There exists a deterministic polynomial-time algorithm $A$ that, given an instance $x\in I$, computes some solution $s\in F_{I}(x)$.
@@
    There exists a deterministic polynomial-time algorithm $B$ that, given $x\in I$ and $s\in F_{I}(x)$, computes the value of $c(s,x)$.
@@
    There exists a deterministic polynomial-time algorithm $C$ that, given $x\in I$ and $s\in F_{I}(x)$, either computes a solution $s^{\prime}\in\mathcal{N}(s,x)$ with $c(s^{\prime},x)>c(s,x)$ (in the case of a maximization problem), or outputs that such a solution does not exist.
@@
Definition 2.
Let $P$ be a local search problem and let $I$ be an instance of P. Moreover, let $s^{*}(I)$ be the unique local optimum obtained by starting with the solution outputted by algorithm $A$ and replacing the current solution by the better solution outputted by $C$, until reaching a local optimum. The standard algorithm problem for $P$ asks for a given instance $I$ of $P$ and a locally optimal solution $s^{\prime}$ for $I$ with respect to $\mathcal{N}$, whether $s^{\prime}$ is exactly the solution $s^{*}(I)$.
It was shown that for many local search problems the standard algorithm problem is PSPACE-complete [14, 37, 38, 42].
@@
    There exist polynomial-time computable functions $f,g$, such that $f$ maps instances $x$ of $P$ to instances $f(x)$ of $Q$, and $g$ maps pairs (solution s of $(\text{solution $s$ of $f(x)$},x)$ to feasible solutions of $P$.
@@
In addition to this standard notion of a PLS-reduction, Schäffer and Yannakakis [42] defined so-called tight reductions. Given PLS problems $P$ and $Q$ and a PLS-reduction $(f,g)$ from $P$ to $Q$, the reduction is called tight if for any instance $x$ of $P$ we can choose a subset $\mathcal{R}$ of the feasible solutions of $f(x)$ of $Q$ such that:
@@
Tight PLS-reductions have two desired properties [1, Chapter 2]. Suppose $P$ reduces to $Q$ via a tight reduction. First, if the standard algorithm problem for $P$ is PSPACE-complete, then the standard algorithm problem for $Q$ is PSPACE-complete as well. Second, if there exists an instance $x$ of $P$ such that there exists a solution of $x$ that is exponentially far away from any local optimum, then such an instance exists for $Q$ as well. Note that this first property holds irrespective of the choices made by the algorithm $C$ for $Q$ [42].
@@ 2.2 Definitions of Local Search Problems
We will be concerned with various local search problems. In the following we provide a summary of the problems that appear in this paper, and provide definitions for each. Some of the problems considered in this paper are not the most natural ones, but we need them as intermediate problems for our reductions. Moreover, these problems might be useful to show PLS-hardness of other problems having cardinality constraints.
@@
The Flip neighborhood as defined above is perhaps the simplest neighborhood structure for a partitioning problem. For each problem in the list below, we consider only the Flip neighborhood in this paper. Recall that the Flip heuristic for $k$-Means is also referred to as the Hartigan–Wong method [21].
@@
Input: A graph $G=(V,E)$ with non-negative edge weights $w:E\to\mathds{Z}_{\geq 0}$.
Output: A partition $(X,Y)$ of $V$ such that $w(X,Y)$ is maximal.
@@
The definitions of Odd Max/Min Bisection are somewhat unconventional, as one usually considers these problem with an even number of vertices and with the Swap neighborhood, where two solutions are neighbors if one can be obtained from the other by swapping a pair of vertices between parts of the partition. Hardness of Max Bisection/Swap was shown by Schäffer and Yannakakis [42] in a short reduction from Max Cut/Flip, providing as a corollary also a simple hardness proof for the Kernighan-Lin heuristic [28]. The reason we require the Flip neighborhood is that we aim to reduce this problem to Squared Euclidean Max Cut/Flip, where we run into trouble when we use the Swap neighborhood (see Section 3 for details).
@@
Input: A complete graph $G=(V,E)$ with non-negative edge weights $w:E\to\mathds{Z}_{\geq 0}$, such that there exists an injection $f:V\to\mathds{R}^{d}$ such that $w(uv)=\|f(u)-f(v)\|^{2}$ for all edges $uv\in E$.
Output: A partition $(X,Y)$ of $V$ such that $w(X,Y)$ is maximal.
Euclidean Max Cut is defined similarly; the only difference is that we have $w(uv)=\|f(u)-f(v)\|$. Squared Euclidean Max Cut is more commonly defined in terms of an input point set $\mathcal{X}\subseteq\mathds{R}^{d}$, with solutions being partitions $(X,Y)$ of $\mathcal{X}$ with cost $\sum_{x\in X}\sum_{y\in Y}\|x-y\|^{2}$. However, in this formulation, membership of Squared Euclidean Max Cut/Flip in PLS is no longer necessarily true111 The problem Sum of Square Roots asks whether for given numbers $a_{i}$ and $k_{i}$ the linear combination $\sum_{i}k_{i}\sqrt{a_{i}}$ evaluates to zero (among other variants). It is not currently known whether this problem can be solved in time polynomial in the number of bits needed to represent the given linear combination, although a randomized algorithm can solve the problem in polynomial time [12]. Since computing the cost of a solution and deciding local optimality for (Squared) Euclidean Max Cut/Flip and $k$-Means/Flip requires the computation of such linear combinations when formulated in the traditional way, it is not clear whether these formulations are in PLS. Nevertheless, our reductions still imply PLS-hardness for these traditional formulations, as well as the existence of certain hard instances. . Hence, we opt for the above graph-theoretical formulation.
@@
Input: A complete graph $G=(V,E)$ with non-negative edge weights $w:E\to\mathds{Z}_{\geq 0}$, such that there exists a injection $f:V\to\mathds{R}^{d}$ such that $w(uv)=\|f(u)-f(v)\|^{2}$ for all edges $uv\in E$.
Output: A partition $(C_{1},\ldots,C_{k})$ of $V$ such that $\sum_{i=1}^{k}\frac{1}{|C_{i}|}\sum_{u,v\in C_{i}}w(uv)$ is minimal.
The sets $\{C_{1},\ldots,C_{k}\}$ are called clusters. As with Squared Euclidean Max Cut, in the literature $k$-Means is usually specified with an input point set $\mathcal{X}\subseteq\mathds{R}^{d}$, and the cost of a partition $\{C_{1},\ldots,C_{k}\}$ of $\mathcal{X}$ is given as $\sum_{i=1}^{k}\sum_{x\in C_{i}}\|x-\operatorname{cm}(C_{i})\|^{2}$, where $\operatorname{cm}(C_{i})$ denotes to centroid of $C_{i}$. This is equivalent to our formulation when the coordinates of the points are algebraic numbers [40, Proposition 1]. In Lemmas 12 and 13 we will also use this notation.
Positive Not-All-Equal $k$-Satisfiability (Pos NAE $k$-SAT)
Input: A boolean formula with clauses of the form $\operatorname{NAE}(x_{1},\ldots,x_{\ell})$ with $\ell\leq k$, where each clause is satisfied if its constituents, all of which are positive, are neither all true nor all false. Each clause $C$ has a weight $w(C)\in\mathds{Z}$.
Output: A truth assignment of the variables such that the sum of the weights of the satisfied clauses is maximized.
In Odd Half Pos NAE $k$-SAT, additionally the number of variables is odd and it is required that the number of true variables and the number of false variables in any solution differ by exactly one. This is analogous to the relationship between Max Cut and Odd Max Bisection.